OpenStudy (ellecullen):

Homework question: Please help asap a right triangle has sides of length x, x-2, and x-9. what is the length of the shortest side?

OpenStudy (kl0723):

x-9

OpenStudy (ellecullen):

Thank you, How would this be solved? :)

OpenStudy (kl0723):

you can give a value to x... lets say x=20 if x=20 one side is 20 side 2 is 20-2=18 and side 3 is 20-9=11

OpenStudy (kl0723):

so the smallest is 11 ;)

jimthompson5910 (jim_thompson5910):

 you need to solve for x a^2 + b^2 = c^2 (x-2)^2+(x-9)^2 = x^2

OpenStudy (ellecullen):

Thank you, I am using the Pythagorean theorem and the other method